# references.bib

### v1

@@ INTRODUCTION
ML models are unfair due to biases in the data, in the model, and/or in the deployment of the model (Figure 1) [rajkomar_ensuring_2018, mehrabi_survey_2021, fazelpour_algorithmic_2021, mhasawade2021machine, wang_bias_2022, xu_algorithmic_2022, drukker_toward_2023, ferrara_fairness_2024]. The data used in healthcare applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [iliadis2016critical, suresh2021framework]. In addition to societal bias, data can also fail to represent the population of interest due the sampling method, time of collection, or data quality issues (e.g., missing, mismeasured, or not enough data). In the fairness literature, data with these undesirable properties are informally referred to as “biased” [chouldechova_fair_2017, barocas_fairness_2023]. Choices made during model training and evaluation can also incorporate and amplify bias in the data, such as selecting an inappropriate outcome, model, or metric(s) and data used for evaluation. During deployment, bias can be introduced when users selectively disregard or overly trust a model’s outputs, which can further reinforce existing societal bias. Table 1 details some common sources of bias together with an example from the healthcare literature. A more comprehensive review of causes of unfairness and applications outside of healthcare can be found in [suresh2021framework] an [mehrabi_survey_2021].
@@
We introduce a small amount of notation to introduce mathematical definitions of various criteria that complement our discussion in the main text. We denote the outcome of interest as $Y$, the features used for model training as ${\bf X}$ [^2] Note that ${\bf X}$ does not have to be a vector; it could be an image, text, tensor, or any other type of data., and the categorical variable for the protected attribute as $A$. ${\bf X}$ may or may not contain $A$ and we discuss this issue when we introduce individual fairness. We let $S=\hat{f}({\bf X})$ be the output from the ML model, where $\hat{f}$ is learned from a set of training data. For example, when $Y$ is continuous, $S$ is simply the predicted value of $Y$. In the context of classification, $Y$ is a binary label for membership in the positive or negative class and $S$ is the predicted probability of being in the positive class, referred to as the risk score. The final prediction, $D$, is positive if $S$ exceeds a pre-determined classification threshold and negative otherwise.
@@
Separation requires that the risk score does not depend on the protected attribute within the positive and negative classes (i.e., $S\perp\!\!\!\perp A\mid Y)$. This implies that, among individuals in the positive (or negative) class, the rate of making a positive (or negative) prediction is consistent across groups. Common separation-based metrics therefore aim to equalize error rates across the groups, including the false negative rate (known as equality of opportunity), false positive rate (known as predictive equality), or both (known as equalized odds). These metrics have been widely used in clinical settings as misclassified predictions can have severe consequences in practice. For instance, a recent study evaluated equality of opportunity in the context of chest X-ray diagnostics, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [seyyed-kalantari_underdiagnosis_2021]. State-of-the-art computer vision models were found to have higher false negative rates in several intersectional under-served subpopulations, such as Hispanic female patients. There are also situations where prioritizing equalized odds is appropriate [hardt_equality_2016]. In the context of rapid screening tools for COVID-19 in emergency departments, a high false negative rate can lead to insufficient monitoring of COVID-19 while a high false positive rate can result in unnecessary additional testing [yang2023algorithmic]. Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [kleinberg2017inherent].
@@
Sufficiency aims to equalize error rates among individuals with similar predictions [castelnovo_clarification_2022]. Formally, sufficiency requires that the label does not depend on the protected attribute given the risk score ( i.e., $Y\perp\!\!\!\perp A\mid S)$. The risk score is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [barocas_fairness_2023]. Common sufficiency-based metrics focus on equalizing the positive predictive value (known as predictive parity), both the positive and negative predictive value (known as conditional use accuracy equality), and calibration (known as well-calibration )[chouldechova_fair_2017, berk_fairness_2018, kleinberg2017inherent, guo_calibration_2017]. Within healthcare, ML models that fail to meet sufficiency criteria have led to inequitable resource allocation. For example, in models predicting in-hospital patient deterioration, achieving predictive parity ensures that the allocation of intensive care unit beds is based solely on medical need rather than protected attributes so that care is provided to those most in need [byrd2023validation]. In terms of calibration-based metrics, a commercial algorithm to enroll US patients into high-risk was shown to violate well-calibration. At equivalent risk scores, Black patients were in significantly poorer health than white patients and therefore had to be sicker for enrollment into the program [obermeyer_dissecting_2019].
@@
Independence ($S\perp\!\!\!\perp A$)
@@
Separation ($S\perp\!\!\!\perp A\mid Y$)
@@
Sufficiency ($Y\perp\!\!\!\perp A\mid S$)
@@
Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all classification thresholds for males and females [roosli_2022_peeking].
Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all classification thresholds [roosli_2022_peeking].
@@
We conclude with a brief overview of causal fairness, a topic that has gained popularity in recent years and itself worth a tutorial in the context of healthcare. As the name suggests, causal fairness criteria evaluate the causal relationship between a protected attribute and a model’s prediction rather than a performance measure or distance metric as in group and individual fairness [plecko_causal_2022]. A deep understanding of causal fairness requires knowledge of causal inference and we refer readers to several general and fairness-focused reviews on the topic [pearl2009causal, hernan2010causal, loftus2018causal, fawkes2022selection, makhlouf2022survey, carey2022causal, nilforoshan_causal_2022]. We introduce one way of incorporating causality into what it means to be fair based on counterfactuals, which are simply “what-if” statements [pearl2009causal, hernan2010causal]. In the context of fairness, the counterfactual is most often the unobserved prediction that would have happened if the protected attribute had been different [kusner_counterfactual_2017, fawkes2022selection, bynum2024new]. One of the first counterfactual-based fairness criteria, simply called counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables are held constant [kusner_counterfactual_2017].
@@
To better explain this concept, we briefly introduce directed acyclic graphs (DAGs). DAGs consist of nodes that represent variables and edges that connect them to illustrate their causal relationships and are used to formulate many causal fairness criteria [pearl2009causal]. For purposes of illustration, consider an over-simplified view of the world in predicting length of stay in the hospital. The DAG in Figure 3 illustrates that the prediction is directly related to previous diagnoses, age, and prescriptions, yet race could indirectly influence the prediction through its impact on access to healthcare, which in turn influences historical diagnoses [pfohl_counterfactual_2019]. When predicting length of stay in the hospital, counterfactual fairness is satisfied if, for example, the predicted length of stay for a black person is the same prediction they would have received if they were counterfactually born a white person, given all other variables in the DAG are held constant. However, counterfactual fairness can generally not be quantified from observed data [kusner_counterfactual_2017, salimi2019interventional, fawkes2022selection].
@@
Other quantities from the causal inference literature have since been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the prediction [pearl2009causality, mitchell_algorithmic_2021, alves_survey_2023]. Conditional counterfactual parity is a weaker criteria that considers the average causal effect of the protected attribute on the prediction conditional on a set of relevant observed variables [kusner_counterfactual_2017]. In our length of stay prediction example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the length of stay in the hospital does not change if their race were counterfactually altered. These definitions, as well as extensions of the previously introduced group and individual fairness measures, are formally introduced in Supplementary Section 6. Further work on causal fairness, such as path-specific and interventional fairness, are detailed in [chiappa_path-specific_2019, nabi_fair_2018, wu_pc-fairness_2019, zhang_causal_2017, salimi2019interventional, zuo_interventional_2024].
@@
Figure 3: Directed acyclic graph (DAG) for predicting length of stay in the hospital. The DAG illustrates the causal relationships among various factors influencing the predicted length of hospital stay. Historical diagnoses, age, and prescriptions have a direct influence on predicted length of stay. Race indirectly affects the prediction through access to care, which in turn influences historical diagnoses.
@@ DISCUSSION
While this paper focuses on how to define fairness, strategies for mitigating bias within ML applications are an equally active area of research. A recent systematic review presents various bias mitigation strategies and their use in practice [hort2023bias]. A necessary step in any bias mitigation strategy is to select the fairness definition that the mitigation strategy will attempt to enforce. Group fairness metrics are the most common targets, though alternative strategies have recently been proposed [goethals2024beyond]. Broadly, bias mitigation can be performed by de-biasing the training data (pre-processing), during the model training (in-processing), and after model training (post-processing). Pre-processing techniques include resampling or reweighting samples to correct imbalances among groups. One approach to in-processing is to include additional regularization terms to penalize a model’s deviation from a pre-specified fairness metric. In post-processing, for example, a technique to achieve equalized odds is to set distinct decision thresholds for different groups [hardt_equality_2016]. However, a practical consideration in any bias mitigation strategy is the well-known trade-off between achieving fairness and high overall accuracy and calibration [menon_cost_2018, brahmbhatt2023towards, mbakwe2023fairness]. Strategies for bias mitigation throughout the ML pipeline have recently been discussed in [gichoya2023ai].
@@
Missing data is a critical challenge to fairness evaluation in healthcare applications. Traditional imputation methods used to address missingness in features have been observed to exacerbate gaps in fairness and fairness-aware imputation strategies have been recently proposed [zhang2021fairness, zhang2022fairness]. Missingness in the outcome as well as the protected attribute can also occur, which brings challenges to both model training and evaluation. For example, in many EHR and medical imaging applications, the outcome of interest is only available for a small subset of data as it is obtained from time-consuming and expensive annotation. Semi-supervised strategies for fair model training and fairness evaluation are an active area of research [wick2019unlocking, ji2020can, choi2020fair, gao2024semisupervised].
In many settings, the protected attribute of interest is may also be unavailable or measured incorrectly. Note that some protected attributes are missing from Table 2, such as religion, gender identity, and sexual orientation, which are rarely available in healthcare data. For instance, most healthcare organizations have yet to integrate sexual orientation data collection into routine care due to the absence of standardized collection practices, the inability of EHR systems to capture structured data, and the socio-cultural challenges of self-reporting [grasso2019planning, kronk_transgender_2022, ruberg2020data]. Advances in fairness have therefore largely omitted sexual orientation, which underscores the necessity of future research to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [tomasev2021fairness].

### v2

@@ INTRODUCTION
ML models are unfair due to biases in the data, in the model, and/or in the deployment of the model (Figure 1) [rajkomar_ensuring_2018, mehrabi_survey_2021, fazelpour_algorithmic_2021, mhasawade2021machine, wang_bias_2022, xu_algorithmic_2022, drukker_toward_2023, ferrara_fairness_2024]. The data used in healthcare applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [iliadis2016critical, suresh2021framework]. In addition to societal bias, data can also fail to represent the population of interest due the sampling method, time of collection, or data quality issues (e.g., missing, mismeasured, or not enough data). In the fairness literature, data with these undesirable properties are informally referred to as “biased” [chouldechova_fair_2017, barocas_fairness_2023]. Choices made during model training and evaluation can also incorporate and amplify bias in the data, such as selecting an inappropriate outcome, model, or metric(s) and data used for evaluation. During deployment, bias can be introduced when users selectively disregard or overly trust a model’s outputs, which can further reinforce existing societal bias. Table 1 details some common sources of bias together with an example from the healthcare literature. A more comprehensive review of causes of unfairness and applications outside of healthcare can be found in [suresh2021framework] and [mehrabi_survey_2021].
@@
We use a small amount of notation to introduce mathematical definitions of various criteria that complement our discussion in the main text. We denote the outcome of interest as $Y$, the features used for model training as ${\bf X}$ [^2] Note that ${\bf X}$ does not have to be a vector; it could be an image, text, tensor, or any other type of data., and the categorical variable for the protected attribute as $A$. ${\bf X}$ may or may not contain $A$ and we discuss this issue when we introduce individual fairness. We let $S=\hat{f}({\bf X})$ be the output from the ML model, where $\hat{f}$ is learned from a set of training data. For example, when $Y$ is continuous, $S$ is simply the predicted value of $Y$. In the context of classification, $Y$ is a binary label for membership in the positive or negative class and $S$ is the predicted probability of being in the positive class, referred to as the risk score. The final prediction, $D$, is positive if $S$ exceeds a pre-determined classification threshold and negative otherwise.
@@
Separation requires that the prediction does not depend on the protected attribute within the positive and negative classes (i.e., $D\perp\!\!\!\perp A\mid Y)$. This implies that, among individuals in the positive (or negative) class, the rate of making a positive (or negative) prediction is consistent across groups. Common separation-based metrics therefore aim to equalize error rates across the groups, including the false negative rate (known as equality of opportunity), false positive rate (known as predictive equality), or both (known as equalized odds). These metrics have been widely used in clinical settings as misclassified predictions can have severe consequences in practice. For instance, a recent study evaluated equality of opportunity in the context of chest X-ray diagnostics, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [seyyed-kalantari_underdiagnosis_2021]. State-of-the-art computer vision models were found to have higher false negative rates in several intersectional under-served subpopulations, such as Hispanic female patients. There are also situations where prioritizing equalized odds is appropriate [hardt_equality_2016]. In the context of rapid screening tools for COVID-19 in emergency departments, a high false negative rate can lead to insufficient monitoring of COVID-19 while a high false positive rate can result in unnecessary additional testing [yang2023algorithmic]. Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [kleinberg2017inherent].
@@
Sufficiency aims to equalize error rates among individuals with similar predictions [castelnovo_clarification_2022]. Formally, sufficiency requires that the label does not depend on the protected attribute given the prediction ( i.e., $Y\perp\!\!\!\perp A\mid D)$. The prediction is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [barocas_fairness_2023]. Common sufficiency-based metrics focus on equalizing the positive predictive value (known as predictive parity), both the positive and negative predictive value (known as conditional use accuracy equality), and calibration (known as well-calibration )[chouldechova_fair_2017, berk_fairness_2018, kleinberg2017inherent, guo_calibration_2017]. Within healthcare, ML models that fail to meet sufficiency criteria have led to inequitable resource allocation. For example, in models predicting in-hospital patient deterioration, achieving predictive parity ensures that the allocation of intensive care unit beds is based solely on medical need rather than protected attributes so that care is provided to those most in need [byrd2023validation]. In terms of calibration-based metrics, a commercial algorithm to enroll patients into a high-risk care management program was shown to violate well-calibration. At equivalent risk scores, Black patients were in significantly poorer health than white patients and therefore had to be sicker for enrollment into the program [obermeyer_dissecting_2019].
@@
Independence ($D\perp\!\!\!\perp A$)
@@
Separation ($D\perp\!\!\!\perp A\mid Y$)
@@
Sufficiency ($Y\perp\!\!\!\perp A\mid D$)
@@
Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all values of the risk score for males and females [roosli_2022_peeking].
Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all values of the risk score [roosli_2022_peeking].
@@
We conclude with a brief overview of causal fairness, a topic that has gained popularity in recent years and itself worth a tutorial in the context of healthcare. As the name suggests, causal fairness criteria evaluate the causal relationship between a protected attribute and a model’s prediction rather than a performance measure or distance metric as in group and individual fairness [plecko_causal_2022]. A deep understanding of causal fairness requires knowledge of causal inference and we refer readers to several general and fairness-focused reviews on the topic [pearl2009causal, hernan2010causal, loftus2018causal, fawkes2022selection, makhlouf2022survey, carey2022causal, nilforoshan_causal_2022]. We introduce one way of incorporating causality into what it means to be fair based on counterfactuals, which are simply “what-if” statements [pearl2009causal, hernan2010causal]. In the context of fairness, the counterfactual can be the unobserved prediction that would have happened if the protected attribute had been different [kusner_counterfactual_2017, fawkes2022selection, bynum2024new]. One of the first counterfactual-based fairness criteria, simply called (individual) counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables are held constant [kusner_counterfactual_2017].
@@
To better explain this concept, we briefly introduce directed acyclic graphs (DAGs). DAGs consist of nodes that represent variables and edges that connect them to illustrate their causal relationships and are used to formulate many causal fairness criteria [pearl2009causal]. For purposes of illustration, consider an over-simplified view of the world in predicting length of stay in the hospital. The DAG in Figure 3 illustrates that the prediction is directly related to previous diagnoses, age, and prescriptions, yet race could indirectly influence the prediction through its impact on access to healthcare, which in turn influences historical diagnoses [pfohl_counterfactual_2019]. When predicting length of stay in the hospital, counterfactual fairness is satisfied if, for example, the predicted length of stay for a black person is the same prediction they would have received if they were counterfactually born a white person, given all other variables are held constant. However, counterfactual fairness can generally not be quantified from observed data [kusner_counterfactual_2017, salimi2019interventional, fawkes2022selection].
@@
Other quantities from the causal inference literature have since been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the prediction [pearl2009causality, mitchell_algorithmic_2021, alves_survey_2023]. Conditional counterfactual fairness is a stronger criteria that considers the total effect of the protected attribute on the prediction conditional on a set of relevant observed variables [kusner_counterfactual_2017]. In our length of stay prediction example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the length of stay in the hospital does not change if their race were counterfactually altered. These definitions, as well as extensions of the previously introduced group and individual fairness measures, are formally introduced in Supplementary Section 6. Further work on causal fairness, such as path-specific and interventional fairness, are detailed in [chiappa_path-specific_2019, nabi_fair_2018, wu_pc-fairness_2019, zhang_causal_2017, salimi2019interventional, zuo_interventional_2024].
@@
Figure 3: Directed acyclic graph (DAG) for predicting length of stay in the hospital. The DAG illustrates the relationships among various factors influencing the predicted length of hospital stay. Historical diagnoses, age, and prescriptions have a direct influence on predicted length of stay. Race indirectly affects the prediction through access to care, which in turn influences historical diagnoses.
@@ DISCUSSION
While this paper focuses on how to define fairness, strategies for mitigating bias within ML applications are an equally active area of research. A recent systematic review presents various bias mitigation strategies and their use in practice [hort2023bias]. A necessary step in any bias mitigation strategy is to select the fairness definition that the mitigation strategy will attempt to enforce. Group fairness metrics are the most common targets, though alternative strategies have recently been proposed [goethals2024beyond]. Broadly, bias mitigation can be performed by de-biasing the training data (pre-processing), during model training (in-processing), and after model training (post-processing). Pre-processing techniques include resampling or reweighting samples to correct imbalances among groups. One approach to in-processing is to include additional regularization terms to penalize a model’s deviation from a pre-specified fairness metric. In post-processing, for example, a technique to achieve equalized odds is to set distinct decision thresholds for different groups [hardt_equality_2016]. However, a practical consideration in any bias mitigation strategy is the well-known trade-off between achieving fairness and high overall accuracy and calibration [menon_cost_2018, brahmbhatt2023towards, mbakwe2023fairness]. Strategies for bias mitigation throughout the ML pipeline have recently been discussed in [gichoya2023ai].
@@
Missing data is a critical challenge to fairness evaluation in healthcare applications. Traditional imputation methods used to address missingness in features have been observed to worsen gaps in fairness and fairness-aware imputation strategies have been recently proposed [zhang2021fairness, zhang2022fairness]. Missingness in the outcome as well as the protected attribute can also occur, which brings challenges to both model training and evaluation. For example, in many EHR and medical imaging applications, the outcome of interest is only available for a small subset of data as it is obtained from time-consuming and expensive annotation. Semi-supervised strategies for fair model training and fairness evaluation are an active area of research [wick2019unlocking, ji2020can, choi2020fair, gao2024semisupervised].
In many settings, the protected attribute of interest may also be unavailable or measured incorrectly. Note that some protected attributes are missing from Table 2, such as religion, gender identity, and sexual orientation, which are rarely available in healthcare data. For instance, most healthcare organizations have yet to integrate sexual orientation data collection into routine care due to the absence of standardized collection practices, the inability of EHR systems to capture structured data, and the socio-cultural challenges of self-reporting [grasso2019planning, kronk_transgender_2022, ruberg2020data]. Advances in fairness have therefore largely omitted sexual orientation, which underscores the necessity of future research to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [tomasev2021fairness].